OpenStudy (anonymous):

an ipod touch is rectangular in shape and has a perimeter of 343.6 mm find the length and width given that it is 48.2mm longer than it is wide

OpenStudy (anonymous):

idk what to do ?

OpenStudy (ganpat):

Perimeter of rectangle = 2 (l + b) and given, wide = x, Therefore l = (x + 48) and Perimeter = 343.6 can u solve further ??

OpenStudy (anonymous):

thank you :)

OpenStudy (ganpat):

wc :)